A metallic rod 1 cm long with a square cross-section is
heated through 1°C. If Young's modulus of elasticity of
the metal is E and the mean coefficient of linear
expansion is alpha per degree Celsius, then the
compressional force required to prevent the rod from
expanding along its length is : (Neglect the change of cross
sectional area)

Grade:12th pass

2 Answers

Vikas TU
4 years ago
Thermal expansion in form of change in length 
l2 = l1(1+alpha*t)
Young’s modulus = E 
E = (F/A)/(l/L)
E = FL/Al
L = l + del l 
E = F (l+del l )/A* del l 
Find the value of F 
F = EA(alpha *t)/(1+ alpha *t)
 
Khimraj
4 years ago
Thermal expansion in form of change in length (Linear expansion) l 2 = l 1 ( 1 + α t ) l2=l1(1+αt) c h a n g e i n l e n g t h = ( l Δ t ) change in length=(lΔt) Δ t = ( t 2 − t 1 ) Δt=(t2−t1) Y o u n g ′ s M o d u l u s = E Young′s Modulus=E E = F A / l L E=FA/lL E = F L / A l E=FL/Al L = ( l + Δ l ) L=(l+Δl) L = l + Δ l L=l+Δl E = F A / l L E=FA/lL E = F L / A l E=FL/Al E = F ( l + Δ l ) A Δ l E=F(l+Δl)AΔl F = E A Δ l / ( l + Δ l ) F=EAΔl/(l+Δl) F = E A ( α t ) 1 + α t F=EA(αt)1+αt